A train starts from rest at a station and accelerate at a rate of 2 m/s2for 10 s. It then runs at constant speed for 30 s, and slows down at -4 m/s2 until it stops at the next station. Find the total distance covered


1
Expert's answer
2021-10-24T18:29:38-0400

At the first part of journey the distance can be found from the kinematic equation:


"d_1 = \dfrac{a_1t_1^2}{2} = \dfrac{2m/s^2\cdot (10s)^2}{2} = 100m"

where "a_1 = 2m\/s^2,t_1 = 10s".

The speed gained by the end of this part is:


"v_2 = a_1t_1 = 2m/s^2\cdot 10s = 20m/s"


At the second part the distance is:


"d_2 = v_2t_2 = 20m/s\cdot 30s = 600m"

And the distance for the last part can be found from another kinematic equation:


"d_3 = \dfrac{v_3^2-v_2^2}{2a_3}"

where "v_3 = 0m\/s" (train stops), "a_3 = -4m\/s^2". Thus, obtain:


"d_3 = \dfrac{-(20m/s)^2}{2\cdot (-4m/s^2)} = 50m"

Finally, the total distance:


"d = d_1 + d_2 + d_3 = 100m + 600m + 50m = 750m"

Answer. 750m.